# The fluctuation-dissipation relation holds for a macroscopic tracer in an active bath

### v1

@@ Abstract
The response of the system to the perturbation arrest, at ${t_{1}}=t-t_{0}$, given by $R_{x}(t_{1})=\delta\langle x(t_{1})\rangle=\langle x(t_{1})\rangle-\langle x\rangle_{0}$, and the unperturbed autocorrelation function, $C_{x}(t_{1})=\langle x(t_{1})x(0)\rangle_{0}$, are therefore measured from parts III, and IV respectively (Fig. 3C). According to the generalized FDR, Eq. 1 should hold exactly for our measured response and autocorrelation if we define $T=T_{\text{eff}}$ and $F_{0}=k\Delta x$, where $\Delta x\equiv\langle x\rangle_{F_{0}}-\langle x\rangle_{0}$. This is verified in Fig. 3D with good accuracy, where we have used the value of $T_{\text{eff}}$ obtained from the stationary position distribution. We note that at $t_{1}>T_{c}=1$ s the stationary noise becomes significant and the agreement or disagreement of $C_{x}$ and $R_{x}$ is hard to assess. Therefore, we cannot rule out an FDR violation beyond $T_{c}$.
Since $T_{\text{eff}}$ fulfills the generalized FDR for any $t_{1}<T_{c}$, it also fulfills it at $t_{1}=0$. Therefore, $k_{B}T_{\text{eff}}=k\langle\Delta x^{2}\rangle_{0}$, which is trivial for a system exhibiting Boltzmann-like statistics. However, this relation should generally hold for any $P(x,y)$ (with different $N_{b}$).
We test the validity of the generalized FDR further by varying the perturbation amplitude (fan voltage) at constant $N_{b}$ and $T_{\text{eff}}$. Figure 4A shows an increase in the measured mean displacements $\Delta x$ as a function of the fan operating voltage. We find that when $\Delta x>\sigma_{x}$ for $V=13.5$ V (Fig. 4A), we are beyond the linear response regime. Consequently, in Fig. 4B we observe that deviations from a generalized FDR become notably more pronounced under stronger perturbations.
Our experimental findings establish a consistent effective temperature, $T_{\text{eff}}$, derived from the generalized fluctuation-dissipation relation (FDR). This $T_{\text{eff}}$ is consistent with the temperature determined by the average potential energy and remains independent of the force. With this defined $T_{\text{eff}}$, we can further explore the conditions under which the generalized FDR remains applicable.
Range of validity of the FDR.– Finally, we repeat our experiments for different numbers of bbots in the arena, $174<N_{b}<$ 17. In Fig. 5A (upper panel) the displayed characteristic times exhibit a dependence on the system configuration $N_{b}$, i.e., the evaluated mean free time between collisions $\tau_{c}$ and the Langevin typical times $\tau_{r}$, $\tau_{\Omega}$ (fit to Eq. 2). We observe that $\tau_{r}$ and $\tau_{c}$ decrease with $N_{b}$, as the dynamics become less oscillatory, approaching a critically-damped behavior at $N_{b}=15,~{}\text{and}~{}17$. By using a consistent definition of effective temperature for all configurations, $T_{\text{eff}}=k\langle\Delta x^{2}\rangle_{0}$, we observe a monotonic increase in $T_{\text{eff}}$ with $N_{b}$ (Fig. 5A, middle panel). We quantify the departure from Boltzmann statistics by calculating the non-Gaussian parameter of the position distribution, $\Gamma=\frac{\langle x^{4}\rangle}{3\langle x^{2}\rangle^{2}}-1$ (see Fig. 5A, lower panel), and find that $P(x,N_{b})$ transitions into a non-Gaussian distribution as $N_{b}$ decreases (Fig. 5B).
The generalized FDRs for different $N_{b}$ are displayed in Fig. 5C and are empirically satisfied within the response time regime. We, therefore, find that a generalized FDR, with an effective temperature as defined here, is generally valid for all our experiments regardless of the functional form of the stationary position distribution, both for underdamped and critically-damped dynamics.
@@
Figure 5: Effective temperatures and FDRs of different bbot configurations. We consider configurations of $N_{b}=\{4,6,10,12,15,17\}$ bbot baths. The results were obtained for an ensemble average of $M=375$ perturbation sequences, with the same tracer $m\approx 1$ g, gravitational stiffness $k\approx 2.8~{}\text{g/s}^{-2}$, and fan operating voltage with $\Delta x(V)<\sigma_{x}$. A.  Different metastatistics vs. number of bbots. Top panel: mean free time between collisions $\tau_{c}$ (magenta, line), Langevin timescales (obtained by a fit to Eq. 2) $\tau_{r}$ (red, squares) and $\tau_{\Omega}$ (blue, crosses), and oscillation period in the potential $\sqrt{m/k}$ (dashed line) Middle panel: effective temperature $T_{\text{eff}}$ (circles, middle panel). Lower panel: non-Gaussian parameter (squares), $\Gamma=\langle x^{4}\rangle/3\langle x^{2}\rangle^{2}-1$. Error-bars are standard deviations. B.  Position probability density for $N_{b}=4$ (diamonds), $10$ (squares), $15$ (circles). Dashed line is the normal (Gaussian) distribution for $N_{b}=15$. C.  Corresponding generalized FDRs (Eq. 1 with $T=T_{\text{eff}}$) for all configurations $N_{b}$. The vertical dashed lines are $T_{c}=1$ s.
Discussion.– In summary, we have experimentally confirmed the validity of a generalized FDR, far from equilibrium, in a large range of conditions. Here, the transient response to a perturbation and the stationary correlations are characterized by a time-dependent relaxation towards a steady state. These quantities exhibit equilibrium-like proportionality via $T_{\text{eff}}$, with dynamics that become less oscillatory as $N_{b}$ rises. The similar values of the evaluated $\tau_{r}$ and $\tau_{c}$, indicate that the number of bath particles $N_{b}$ governs both the damping rate and the effective temperature ($T_{\text{eff}}\sim\langle\Delta x^{2}\rangle_{0}$) of the tracer. In essence, the bots play a central role in both energy dissipation and fluctuation generation in the system.
We focused on experimental setups with $174<N_{b}<$ 17 for two main reasons. When $N_{b}<4$, the tracer is rarely perturbed, and conversely, when $N_{b}>17$, finite system size leads to large clusters at the trap’s center. It is noteworthy that although possible violations are observed at increasingly lower times as $N_{b}$ decreases, and the analogy to a heat bath becomes limited, the dynamic FDR with $T_{\text{eff}}\sim\langle\Delta x^{2}\rangle_{0}$ holds, within statistical accuracy ($t_{1}<T_{c}$), even at $N_{b}<12$ when the tracer’s position distribution is highly non-Gaussian.
The confirmation of a generalized FDR, across an extensive range of effective temperatures, raises the following three broader issues. Is there an overarching equipartition relation that connects these effective temperatures to $T_{\text{eff}}$ defined via momentum variables rather than positional ones?
Secondly, do these temperatures hold significance within the thermodynamic framework governing these systems? For instance, could they dictate energy flow in scenarios where two such systems are in contact?
Lastly, our experimental observations reveal that our tracer dynamics exhibit consistency with Markovian dynamics [17]. While Markovianity is not essential for the FDR to prevail in thermal equilibrium, an intriguing avenue of exploration involves assessing its validity in weakly perturbed systems immersed in an active, non-Markovian environment, where dissipation and fluctuation derive from the same underlying physical process.

### v2

@@ Acknowledgements
Figure 12: Passive response measurement. We probe the tracer’s coupling to the trap in the absence of collisions by measuring the (deterministic) response $\delta x_{0}$ to a step-perturbation with the bbots removed ($13.5$V fan voltage, abruptly switched off at $t=0$). Considering the tracer relaxes as a damped harmonic oscillator with natural frequency $(\tau^{0}_{\Omega})^{-1}=\sqrt{k/m}=5.3\ \textrm{s}^{-1}$, the passive damping timescale is obtained by fitting the normalized response to Eq. 2, $\tau^{0}_{r}=0.9$s.
Figure 13: Tracer’s coefficient of restitution versus impact velocities. The restitution coefficient $r$ was determined from the sound signal emitted by the styrofoam ball (the tracer) bouncing repeatedly off a surface, as detailed in Ref. [61]. We obtain $r$ by measuring the time lag between consecutive impacts: $r_{i}=|v^{\prime}_{i}/v_{i}|=(t_{i+1}-t_{i})/(t_{i}-t_{i-1})$, for an impact velocity $v_{i}=\frac{g}{2}(t_{i-1}-t_{i})$, $i=2,3,...$. Here, $v_{i}$ and $v^{\prime}_{i}$ are the normal pre- and post-collision velocities occurring at time $t_{i}$, and $g=9.81\textrm{m/s}^{2}$ is the gravitational acceleration. The plot shows four successive impacts of a repeated experiment (50 trials), with mean values from last to first respectively: $r=0.79$ (circle), $0.78$ (diamond), $0.76$ (square), $0.74$ (triangle), for $v_{\perp}=68.37$, $87.35$, $113.4$, $152.83$ cm/s.
Figure 14: Kinetic and potential temperatures. The potential temperature, $k_{B}T_{x}=k\langle\Delta x^{2}\rangle_{0}$, is plotted versus the kinetic temperature of the tracer, $k_{B}T_{v}=m\langle v^{2}\rangle_{0}$. We find a linear relation for the range of $3\geq N_{b}\leq 10$. The red line is the best linear fit for this range.
@@
Figure 15: FDR results with different numbers of bbots.  A,B. Generalized FDRs with $T_{\text{eff}}\sim\langle\Delta x^{2}\rangle_{0}$ for $N_{b}=\{2,3,4,6,10,12,15,17\}$ bbot baths. The results were obtained for an ensemble average of $M=375$ perturbation sequences, with the same tracer $m\approx 1$ g, gravitational stiffness $k\approx 28.2~{}\text{g/s}^{-2}$, and fan operating voltage with $\Delta x(V)<\sigma_{x}$. The vertical dashed lines are $T_{c}=1$ s.